OpenStudy (anonymous):

use the double angle identities to rewrite sin^2(x) in terms of linear powers of common trig functions

OpenStudy (dumbcow):

\[\cos(2x) = 1-2\sin^{2} x\] work backwards

OpenStudy (anonymous):

@dumbcow where did the 2 in front of the sin come from?

OpenStudy (dumbcow):

\[\cos(2x) = \cos^{2} x - \sin^{2} x = (1-\sin^{2} x) -\sin^{2} x = 1-2\sin^{2} x\]

OpenStudy (anonymous):

@dumbcow ok but how did you get the cos(2x) from just sin^2 x?

OpenStudy (dumbcow):

trig identities :) ....get a reference sheet, there a lot of trig identities to remember look up double angle identities for cosine and the 1 i posted will come up

OpenStudy (anonymous):

is it sin^2x=(1-cos2x)/2

OpenStudy (dumbcow):

yep

OpenStudy (dumbcow):

if you want to know where the identity came from ....look up angle sum formula for cos cos(2x) = cos(x+x) = cos^2 x - sin^2 x

OpenStudy (anonymous):

ok thanks! what does it mean by "in terms of linear powers of common trig functions" how will i know when ive found the answer?

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

oh thats fancy language meaning answer only has sin/cos/tan with no powers

OpenStudy (anonymous):

ok does the cos2x count?

OpenStudy (dumbcow):

yes because its cos to first power...doesn't matter if its 2x or 3x inside parenthesis

OpenStudy (anonymous):

ok! im stuck after cos2x=1-sin^2x i dont know how to get rid of the sin^2x

OpenStudy (dumbcow):

?? you already did it \[\sin^{2} x = \frac{1-\cos(2x)}{2}\]

OpenStudy (anonymous):

oh thats all it was asking?!

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

oh lol